OpenStudy (joshoyen):

Write the slope-intercept form of the equation of each line. 1. through (4 2) perp. to y = -2/3x -4

OpenStudy (anonymous):

do you know the slope of \[\huge y=\color{red}{-\frac{2}{3}}x-4\]

OpenStudy (joshoyen):

the slope is -2/3 ?

OpenStudy (anonymous):

that is the slope of that line, yes how about the slope of a perpendicular line?

OpenStudy (joshoyen):

is it the opposite of it? so +2/3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not the "opposite" although you do change the sign it is the "negative reciprocal" so flip it and change the sign

OpenStudy (joshoyen):

so it is 3/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you good from there?

OpenStudy (joshoyen):

and then you do the y-y1=m(x-x1) formula right?

OpenStudy (anonymous):

zactly

OpenStudy (joshoyen):

aight thanks a lot!

OpenStudy (anonymous):

yw